Solve Linear Equation of f(x) = g(x)
Given:
f(x) = 2x + 5
g(x) = 5x/2

Answer by rapaljer(4671)

f(x)=g(x)
2x+5=2.5x

Subtract 2x from each side:
5=.5x

Multiply both sides by 2:
2*5=2*.5x
10=x

Dr. Robert J. Rapalje, Retired
Professor of Mathematics
Seminole State College of Florida